A ship called at an African port, but was too large to tie up at a whatf. The whole cargo was unloaded by 25 barges, each taking 3 loads. If each barge had been large enough to hold 160 tonnes more, there would have been only one load each. How heavy was the ships cargo?
Answer by Theo(13342)

let x = the number of tons of cargo that each barge can carry in one load.

each barge carried 3 loads each.

the total tons of cargo carried was therefore 3 * 25 * x which is equal to 75 * x.

if each barge had been able to carry 160 more tons per load, then the 25 barges would only have had to carry 1 load apiece.

the total tons of cargo carried would therefore have been 1 * 25 * (x + 160) which is equal to 25 * (x + 160).

since the total tons of cargo that was unloaded had to be the same, you get:

75 * x = 25 * (x + 160).

simplify to get 75 * x = 25 * x + 4000

subtract 25 * x from both sides of the equation to get 50 * x = 4000

divide both sides of the equation by 50 to get x = 80.

each barge was able to carry 80 tons of cargo in each load.

the total tons of cargo carried was therefore 25 * 3 * 80 = 6000 tons.

if each barge were able to carry 160 more tons each, then each barge would have been able to carry 80 + 160 = 240 tons of cargo each.

they would then have only had to take 1 load each.

25 * 1 * 240 = 6000 total tons of cargo.

either way, 6000 tons of cargo would have been unloaded.

the ship's cargo therefore had to weigh 6000 tons.